# Communication Optimization for Distributed Training: Architecture, Advances, and Opportunities

### v1

@@ III-A Parallelization Strategy
The communication overhead in distributed training mainly arises from parallelization. Different implementations of parallelization strategies affect the communication pattern and network traffic during the model training process [2].
@@
The most straightforward and most widely used parallelization strategy is data parallelism. As mentioned in Section II, multiple copies of the same model are replicated across different nodes in data parallelism. Each iteration requires all nodes to synchronize gradients, which are used for global parameter updates. Consequently, data parallelism results in a significant amount of All-Reduce communication.
@@
To train a large-scale model that can’t fit into the memory of a single GPU, researchers introduced model parallelism, which includes pipeline and tensor parallelism. Tensor parallelism divides models within a layer and requires intensive communication in training. Researchers from Nvidia proposed an efficient “intra-layer parallelism” approach called Megatron-lm [7] to reduce the communication traffic in tensor parallelism. Pipeline parallelism necessitates only the transfer of activations/gradients between adjacent nodes in the logical pipeline, which is a point-to-point communication pattern that significantly reduces the communication cost compared to data parallelism and tensor parallelism. Because overlapping communication and computation is hard to achieve in naive pipeline parallelism, an interleaved pipeline scheduling strategy [1] is proposed where each device is assigned multiple data chunks to gain the chance of communication-computation overlap. However, pipeline parallelism, as an “inter-layer parallelism” approach, is not without its flaws. It faces issues such as “synchronous/asynchronous” parameter updates and pipeline bubble problems, which can adversely affect the accuracy of the model and the efficiency of training [7].
@@
Various parallelization strategies are often used in conjunction. For example, PTD-P [1] not only combines Megatron-lm’s Tensor parallelism and pipeline parallelism but also introduces data parallelism to scale to thousands of GPUs. AlpaComm [8] identifies the “cross-mesh resharding” problem that arises when combining tensor and pipeline parallelism and proposes a communication optimization method based on broadcasting and an “overlapping-friendly” pipeline scheduling scheme to accelerate end-to-end training.
@@
As an emerging model structure, the MoE model is gradually gaining popularity due to its fast training/inference speed. The training process of the MoE model selects several appropriate experts for a data sample, which inherently has good compatibility with parallelization. MoE parallelism can be achieved by simply dispersing the “experts” on different GPUs, introducing All-to-All traffic when data is distributed between different GPUs. Additionally, the model parameters of the non-expert part are globally shared in MoE parallelism, so there will also be All-Reduce traffic. Lina [9] uses a scheduling strategy that prioritizes All-to-All traffic and further splits the All-Reduce tasks to overlap with computation as much as possible, thus reducing exposed communication and achieving training acceleration. Janus [10] studies MoE parallelism acceleration from a new perspective. Since the size of expert models and training datasets varies, Janus innovatively proposes a “data-centric” model that “moves experts instead of data”. This significantly reduces the amount of communication during the training process under certain conditions (e.g., when the scale of experts is smaller than the scale of data), thus accelerating the training process.
@@ III-B Collective Communication Library
Communication tasks of distributed training are often implemented by CCL. The current predominant NVIDIA Collective Communication Library (NCCL)777https://developer.nvidia.com/nccl dynamically selects established primitive algorithms based on different situations. For instance, NCCL selects the Ring and Double Binary Tree algorithms for All-Reduce 888https://developer.nvidia.com/blog/massively-scale-deep-learning-training-nccl-2-4 for different workloads and topologies to accelerate the training process. However, NCCL’s optimization for specific hardware underscores the need for a more universal and adaptable CCL to speed up distributed training across different hardware and topologies. Therefore, many researchers took different approaches to implement communication primitives more efficiently.
For example, Blink [11] dynamically generates optimal communication primitives by packing spanning trees rather than writing and optimizing them manually. Blink uses integer linear programming (ILP) to reduce the number of trees and improve performance, and also leverages heterogeneous communication channels to utilize full bandwidth. At the time, Blink achieved up to 40% speedup compared to NCCL in end-to-end DNN training iterations.
SCCL [12] takes a more unified approach to achieve efficiency: automatically synthesizing high-performance communication primitives for a given topology. To achieve this, SCCL designs a cost model to evaluate the latency and bandwidth cost of algorithms and then searches for algorithms on the Pareto frontier. However, SCCL also faces problems of high complexity. The communication primitive algorithm generation problem is encoded into a Mixed Integer Linear Program (MILP), which is unfortunately NP-hard.
To reduce the complexity of generating collective communication algorithms, TACCL [5] is a representative work that introduces a “human-in-the-loop” approach, whose paradigm is shown in Fig. 4. TACCL incorporates high-level inputs from an algorithm designer, such as logical topologies, switch hyper-edges, and algorithm symmetry, to efficiently synthesize collective communication algorithms for heterogeneous topologies. These human inputs greatly constrain the search space of algorithms, reducing search time to an acceptable amount, in most cases from seconds to a few minutes. Compared with NCCL, TACCL achieves up to 2.36x end-to-end training speedup for BERT and 1.94x for Transformer-XL.
Different from previous methods focused on speeding up single primitives, SYNDICATE [13] proposes a novel abstraction to overlap primitives in a fine-grained pattern, breaking large communication operations into smaller units, thus allowing higher flexibility. SYNDICATE then develops a joint optimizer based on Markov Chain Monte Carlo (MCMC) search to optimize the joint action of the control plane and the data plane.
@@ III-C Network
Distributed training focuses on GPU-to-GPU communication, which needs to consider intra/inter-host communication protocols. Common intra-host communication protocols include the general PCIe protocol and Nvidia’s NVLink/NVSwitch. Inter-host communication protocols include traditional TCP/IP and more efficient RDMA (e.g., RoCE and Infiniband).
@@
Based on certain communication protocols, the underlying network topology also significantly influences the communication efficiency of distributed training. For instance, as mentioned in Section II, Nvidia utilizes a Ring and fully connected Mesh topology in their DXG series, and Google employs a 3D-Torus topology in their TPUv4 [4]. These topologies contain a large number of ring structures, which enhance the efficiency of Ring-based All-Reduce and Point-to-Point communications (such as pipeline parallelism or broadcast-based hybrid parallelism [8]). However, static topologies can rarely adapt to dynamically changing job requirements, leading researchers to propose TopoOpt [2]. TopoOpt takes advantage of the reconfigurability of optical switches, enabling a scheme that co-optimizes the topology and parallelization strategy. Nevertheless, current optical switching technology still has issues of high reconfiguration latency, making it challenging to perform dynamic topology adjustments between iterations. As such, TopoOpt remains a relatively coarse-grained optimization, that is, reconfiguring the interconnection between training servers of each job only before the job starts and keeps the topology until the training is complete.
@@
Anyway, research on communication optimization for distributed training from a network perspective is still insufficient, which is prospected in Section IV.
III-D Collaboration Design in Current Advances
@@
As observed in Table I, the research advances mentioned above are primarily limited to optimizations within a layer, with less cross-layer co-design. Current advances involving cross-layer optimization can be categorized into two types. The first type is passive collaborative design. For instance, AlpaComm [8] and Lina [9] optimize communication for a specific parallelization strategy, necessitating modifications to communication primitives. Another example is that generative CCL (such as Blink [11], SCCL [12] and TACCL [5]) has to be aware of network topologies because their main significances are to customize collective communication algorithms for specific topologies. The second type of work involves proactive attempts at collaborative optimization yet remains exploratory. For example, SYNDICATE [13] jointly optimizes the implementation and scheduling of communication primitives, allowing parallel execution of primitives. TopoOpt [2] collaboratively optimizes network topology and parallelization strategies, obtaining optimized combination strategies through alternating optimization techniques.
@@
In summary, collaborative communication optimization in distributed training is still a domain that needs further exploration. Section IV details the importance of co-design and potential research opportunities.
@@ IV Opportunities
As noted in Section III, collaborative design has not received widespread attention in current research advances. However, different from general tasks in data centers, the communication traffic patterns in distributed training are known in most cases [6, 14], providing motivations and opportunities for cross-layer collaboration. Furthermore, the optimization objectives for communication tasks are no longer the general-purpose flow complete time (FCT) but rather the job complete time (JCT). In this context, communication tasks are merely a component of the entire distributed training process. Optimizing JCT necessitates considering the dependency between communication and computation instead of just optimizing each communication task. Communication tasks should be optimized to minimize computation blocking time, which can only be achieved through “cross-layer” scheduling to meet this new optimization objective.
To this end, we advocate a communication-efficient five-layer paradigm with (logical) schedulers as the middleware, as illustrated in Fig. 5. The schedulers not only allow lower-layer facilities to perceive the requirements of higher-layer applications but also enable higher-layer applications to be aware of the capabilities of lower-layer facilities. We refer to this as “Vertical” co-design. The schedulers can also perform global scheduling across multiple jobs, achieving “Horizontal” collaborative optimization. Considering the heterogeneous topologies of distributed training and the development of emerging technologies such as in-network aggregation, we also provide a prospect for “Intra-Inter” and “Host-Net” co-design.
@@
“Vertical” co-design: Refers to exchanging necessary information between layers to optimize communication efficiency. For example, CCL can acquire global traffic demand from the parallelization strategy and generate the optimal communication primitive algorithms in conjunction with the network state. Furthermore, cross-layer schedulers can be constructed to strategically schedule communication tasks based on traffic load or optimize task graphs according to the traffic patterns of relevant communication primitives. Echelon [14] is a flow scheduling work that embodies “Vertical” co-design, where novel communication optimization objectives are designed based on the dependency between communication and computation under various parallelization strategies, aiming to optimize communication towards minimizing JCT.
Another example of “Vertical” co-design is that network can also use traffic demand for optimization in distributed training. As the general infrastructure, the network is often unaware of traffic demand. This is natural in common data center scenarios, where the network can only implement passive adjustments such as RTT-based congestion control due to the random nature of flow arrivals. However, traffic patterns are often predictable in distributed training scenarios[6]. Consequently, the network itself can also proactively adjust by taking traffic demands of high-layer training tasks into account, thus enhancing goodput. This in turn reduces communication time and accelerates the distributed training process.
@@
“Horizontal” co-design: Primarily focuses on the coordination among different training jobs or various communication tasks. Within a single training job, the traffic from multiple concurrent collective communication primitives competes for network resources. Similarly, in the multi-job scenario, the traffic from numerous concurrent training jobs also competes for network resources. Consequently, optimization strategies tailored to a single communication primitive or training job may suffer varying degrees of inefficacy. Therefore, it is imperative to employ horizontal co-design at different granularities to fully utilize network resources, ultimately reducing JCT of training tasks. CASSINI [6] attempts horizontal co-design through the perspective of multi-task scheduling. It analyses the consequence of non-coordinated job scheduling in multi-job scenarios, highlighting the potential traffic conflicts arising from the periodic patterns of network resource utilization by different training jobs. Subsequently, CASSINI introduces a design for “staggering peak” scheduling for different jobs, achieving high utilization of network resources, thus enhancing the performance of multiple training jobs.
@@
“Intra-Inter” co-design: Refers to the coordinated utilization of multidimensional network resources, primarily including the “intra-host” and “inter-host” network. The current networks used in distributed training exhibit substantial heterogeneity. The interconnection patterns between different GPUs within the same host include PCIe or NVLink, while those between different hosts include RoCE or Infiniband, etc. Given that GPU (rather than host) is the minimum unit in the communication tasks of distributed training, two modes of GPU-to-GPU communication exist: within the same host and across different hosts. In this context, the issue of heterogeneous link bandwidths within and between hosts is inevitably confronted. An example of “Intra-Inter” co-design is that TACCL [5] utilizes a profiling approach to perceive the characteristics of heterogeneous links, which serves as a basis for optimizing communication primitive algorithms.
@@
“Host-Net” co-design: Primarily refers to the collaboration of computing resources between the end-host and programmable switches within a network. In recent years, with the development of programmable switch technology, in-network aggregation has gradually become a new computing paradigm. In-network aggregation not only fully utilizes the computing capabilities of programmable switches within the network but also reduces the traffic running in the network, achieving multiple purposes in one stroke. However, since existing protocols are designed for end-to-end communication, implementing in-network aggregation usually requires the support of new protocols. For instance, ATP [15] is an in-network aggregation transmission protocol designed for multi-tenant scenarios, which can make full use of the computing capabilities of multi-level programmable switches and degrade to host aggregation when the in-network computing resources are exhausted. This approach also reduces the network traffic for a collective communication task. ATP embodies the concept of “Host-Net” co-design and represents an exemplary work of in-network computing in the field of distributed training. Nevertheless, in-network aggregation has not been widely applied in the production environment due to its complexity. “Host-Net” co-design remains a direction worthy of long-term research.

### v2

@@ III-A Parallelization Strategy
The communication overhead of distributed training stems from the data synchronization requirements brought about by parallelization. Different parallelization strategies determine different communication types, which in turn bring about different network traffic patterns [2].
@@
The most straightforward and most widely used parallelization strategy is data parallelism. As mentioned in Section II, multiple copies of the same model are replicated across different nodes in data parallelism. Each iteration requires all nodes to synchronize gradients, which are used to update global parameters. Consequently, data parallelism results in a significant amount of All-Reduce communication.
However, due to memory constraints, very large models cannot be trained using only data parallelism. To address this, researchers introduced model parallelism, which includes pipeline and tensor parallelism. Tensor parallelism divides a layer of a model, requiring intensive communication for synchronizing activations/gradients at each layer during propagation. To reduce the communication traffic in tensor parallelism, researchers from Nvidia proposed an efficient “intra-layer parallelism” approach called Megatron-lm [7]. It achieved 77% of linear scaling with a 8.3 billion parameters model on 8 GPUs, and 74% of linear scaling on 512 GPUs.
@@
To further divide the model by layers, researchers introduced pipeline parallelism. Pipeline parallelism only requires transferring activations/gradients between adjacent nodes in the logical pipeline, which is a point-to-point communication mode that requires less communication compared to data parallelism and tensor parallelism. Since communication and computation are difficult to overlap in the original pipeline, researchers introduced an interleaved pipeline scheduling strategy [1] where each device is assigned multiple data chunks at a time to gain the chance of communication-computation overlap. However, pipeline parallelism, as an “inter-layer parallelism” approach, is not without its flaws. It faces issues such as “synchronous/asynchronous” parameter updates and pipeline bubble problems, which can adversely affect the accuracy of the model and the efficiency of training [7].
@@
Various parallelization strategies are often used in conjunction. For example, PTD-P [1] combines tensor parallelism, pipeline parallelism and data parallelism to scale to thousands of GPUs. It can train a model with 1 trillion parameters at 502 petaFLOP/s on 3072 GPUs (per-GPU throughput of 52% of theoretical peak). AlpaComm [8] identifies the “cross-mesh resharding” problem that arises when combining tensor and pipeline parallelism. It proposes a communication optimization method based on broadcasting and an “overlapping-friendly” pipeline scheduling scheme to accelerate end-to-end training. It shows a 1.1x speedup compared to their state-of-the-art baseline because of overlapping.
@@
The MoE model is an emerging model structure that is gaining popularity due to its high performance, flexibility and fast training/inference speed. During the training process, each data sample is assigned to several appropriate experts, which inherently has good compatibility with parallelization. MoE parallelism can be achieved by simply dispersing the “experts” on different GPUs. This introduces All-to-All traffic when data is distributed between different GPUs. Additionally, non-expert part of the model will also introduce All-Reduce traffic during training. There are many studies on optimizing MoE parallelism. Lina [9] uses a scheduling strategy that prioritizes All-to-All traffic and further splits the All-Reduce tasks to overlap with computation as much as possible, thus reducing exposed communication. Janus [10] studies MoE parallelism acceleration from a new perspective. Since the size of expert models and training datasets varies, Janus innovatively proposes a “data-centric” model that “moves experts instead of data”. This significantly reduces the amount of communication traffic up to 16x under certain conditions (e.g., when the scale of experts is smaller than the scale of data) and achieves up to 2.06× speedup compared with MoE training system at the time.
@@ III-B Collective Communication Library
Communication tasks of distributed training are often implemented by CCL. The current predominant NVIDIA Collective Communication Library (NCCL)777https://developer.nvidia.com/nccl dynamically selects established algorithms based on different situations. However, NCCL’s optimization for specific hardware underscores the need for a more universal and adaptable CCL to speed up distributed training across different hardware and topologies. Therefore, many researchers take different approaches to implement communication primitives more efficiently.
For example, Blink [11] dynamically generates optimal communication primitives by packing spanning trees rather than writing and optimizing them manually. It uses integer linear programming (ILP) to reduce the number of trees and improve performance, and also leverages heterogeneous communication channels to fully utilize bandwidth. At the time, Blink achieved up to 1.4x speedup compared to NCCL in end-to-end DNN training iterations. SCCL [12] takes a more unified approach to automatically synthesize high-performance communication primitives for a given topology. It designs a cost model to evaluate the latency and bandwidth cost of algorithms and then searches for algorithms on the Pareto frontier. SCCL is 1.14x to 2.2x faster than NCCL’s All-Gather on different sizes of data at the time. However, generating communication primitives is a Mixed Integer Linear Program (MILP) problem, which is unfortunately NP-hard and can require a long time to solve.
Although Blink and SCCL focus on optimization within small heterogeneous topologies and might not be practical to scale to large-cluster scenarios with tens of thousands of GPUs because of complexity, they lay the groundwork for future CCL optimization works.
Based on SCCL, TACCL [5] introduces a “human-in-the-loop” approach to reduce the complexity of generating collective communication algorithms. The workflow of TACCL is shown in Fig. 4. TACCL incorporates high-level inputs from an algorithm designer, such as logical topologies, switch hyper-edges, and algorithm symmetry, to efficiently synthesize collective communication algorithms for heterogeneous topologies. These human inputs greatly constrain the search space of algorithms, reducing search time to an acceptable amount, in most cases from seconds to a few minutes. Compared with NCCL, TACCL achieves up to 2.36x end-to-end training speedup for BERT and 1.94x for Transformer-XL.
Different from previous methods focused on speeding up single primitives, SYNDICATE [13] proposes a novel abstraction to overlap primitives in a fine-grained pattern. It breaks large communication operations into smaller units, thus allowing higher scheduling flexibility. SYNDICATE jointly optimizes the implementation and scheduling of communication primitives, enabling parallel execution of primitives. It improves the speed of training state-of-the-art large models by 21-45%. SYNDICATE embodies the idea of cross-layer collaborative design, but is still a preliminary exploration.
@@ III-C Network
Distributed training focuses on GPU-to-GPU communication, which needs to consider intra/inter-host communication protocols. Common intra-host communication protocols include the general PCIe protocol and Nvidia’s NVLink/NVSwitch, while inter-host protocols include traditional TCP/IP and more efficient RDMA (e.g., RoCE and Infiniband).
@@
Besides protocols, the underlying network topology also significantly influences the communication efficiency of distributed training. For instance, as mentioned in Section II, Nvidia utilizes a Ring and fully-connected Mesh topology in their DGX series, and Google employs a 3D-Torus topology with optical circuit switches (OCSs) in their TPUv4 [4]. These topologies contain a large number of ring structures, which enhance the efficiency of Ring-based All-Reduce and Point-to-Point communications (such as pipeline parallelism or broadcast-based hybrid parallelism [8]).
However, static topologies can rarely adapt to dynamically changing job requirements, motivating researchers to propose TopoOpt [2]. TopoOpt takes advantage of the reconfigurability of optical switches, enabling a scheme that co-optimizes the topology and parallelization strategy. Nevertheless, current optical switching technology still has issues of high reconfiguration latency, making it challenging to perform dynamic topology adjustments between iterations. As such, TopoOpt remains a relatively coarse-grained optimization. It reconfigures the topology only before the job starts and maintains it until the training is complete.
@@
In conclusion, research on communication optimization for distributed training from a network perspective is still insufficient, which requires further exploration.
III-D Collaborative Design in Current Advances
@@
Current advances involving cross-layer optimization can be categorized into two types. The first is passive collaborative design. For instance, AlpaComm [8] and Lina [9] optimize communication for a specific parallelization strategy, necessitating modifications to communication primitives. Another example is that generative CCL (such as Blink [11], SCCL [12] and TACCL [5]) has to be aware of network topologies because their main focuses are to customize collective communication algorithms for specific topologies. The second type is proactive collaborative design. For example, SYNDICATE [13] jointly optimizes the implementation and scheduling of communication primitives, allowing parallel execution of primitives. TopoOpt [2] collaboratively optimizes network topology and parallelization strategies, obtaining optimized combination strategies through alternating optimization techniques. However, these works are still preliminary attempts and require further research.
@@
In summary, collaborative communication optimization in distributed training is still a domain that needs further exploration. Section IV details the importance of collaborative design and potential research opportunities.
@@ IV Opportunities
Different from general communication tasks in data centers, the data traffic patterns of distributed training are predictable in most cases [6, 14]. Furthermore, the optimization objectives of communication tasks are no longer the general-purpose flow completion time (FCT) but rather the job completion time (JCT). Therefore, optimizing JCT necessitates considering the dependency between communication and computation tasks and the relationship between different communication tasks instead of just optimizing a single communication task. These provide motivations and opportunities for collaborative design to optimize efficiency.
In this section, we use a simple case study (Fig. 5(b)) to illustrate the necessity of collaborative design in distributed training. We provide research opportunities from two aspects: the communication-efficient five-layer paradigm and the collaboration of heterogeneous resources.
IV-A Communication-Efficient Five-Layer Paradigm
In Section II, we have introduced the current three-layer paradigm of communication optimization for distributed training, which includes Parallelization strategy, CCL, and Network. In order to promote cross-layer collaboration, we extend the three-layer communication paradigm to a communication-efficient five-layer paradigm, as illustrated in Fig. 5(a), where the red arrows indicate possible information exchange for collaborative optimization.
In the new paradigm, the Parallelization Strategy (Para.), CCL, and Network (Net.) are still the three main layers, while the (logical) schedulers are added as the middleware. The task scheduler is to schedule the communication tasks required by the parallelization strategy, and the flow scheduler is to schedule the communication flows generated by CCL. Standardized interfaces and protocols are required to facilitate seamless communication between layers, which allow lower-layer facilities to perceive the requirements of higher-layer applications and enable higher-layer applications to be aware of the capabilities of lower-layer facilities. Each layer can use shared information to optimize with a broader perspective, achieving “Vertical” co-design.
@@ IV-A Communication-Efficient Five-Layer Paradigm
For example, CCL can generate task-specific communication primitive algorithms based on the global traffic demand from the parallelization strategy and the network state, thus minimizing JCT. Furthermore, cross-layer schedulers can be constructed to strategically schedule communication tasks based on traffic load or optimize task graphs according to the traffic patterns of relevant communication primitives. In the case study of Fig. 5(b), two concurrent flows of Job1 compete for ToR switch at ①, which may lead to congestion. However, the “deadline” of the two flows may be different, and the scheduler can adjust the priority of the flows to avoid blocking the subsequent computation process. In other words, the scheduler can consider both the task graphs and the network constraints to optimize the communication tasks, which is an example of “Vertical” co-design. This collaborative schedule can minimize JCT for the entire Job1, not only FCT for a certain flow. Echelon [14] is a flow scheduling example that embodies “Vertical” co-design, where novel communication optimization objectives are designed based on the dependency between communication and computation under various parallelization strategies, aiming to optimize communication towards minimizing JCT.
@@
The schedulers can also perform global scheduling across multiple jobs, achieving “Horizontal” co-design. In multi-job scenarios, the traffic from numerous concurrent training jobs competes for network resources. As shown in Fig. 5(b), flows of Job1 and Job2 compete for the same ToR switch at ②. Consequently, optimization strategies tailored to a single training job may suffer varying degrees of inefficacy. Therefore, it is imperative to employ horizontal co-design at different granularities to fully utilize network resources, ultimately reducing JCT of training tasks. A potential solution is to implement dynamic resource manager that allocates resources based on real-time demands. CASSINI [6] attempts horizontal co-design through the perspective of multi-task scheduling. It performs “staggering peak” scheduling for different jobs, which improves the utilization of network resources. In addition, horizontal collaboration also enables multi-job Service Level Agreement (SLA) management.
IV-B Collaborate Design for heterogeneous resources
Considering the heterogeneous network topologies of distributed training and the development of emerging technologies such as in-network aggregation, we provide a prospect for “Intra-Inter” and “Host-Net” co-design.
@@ IV-B Collaborate Design for heterogeneous resources
“Intra-Inter” co-design refers to the coordinated utilization of multidimensional network resources, primarily including the “intra-host” and “inter-host” network. GPUs are typically interconnected via fast PCIe or NVLink within a host, while inter-host connections are established through relatively slower RoCE or InfiniBand. In this context, the issue of heterogeneous link bandwidths within and between hosts is inevitably confronted. Let us suppose that Flow1 of Job1 is to implement a Ring All-Reduce among GPUs across two hosts in Fig. 5(b). Then the bottleneck of this communication task will be the bandwidth of the inter-host link, which is slower than the intra-host link. Solving this problem requires real-time sensing and scheduling of heterogeneous link resources. TACCL [5] utilizes a profiling approach to perceive the characteristics of heterogeneous links in order to optimize the communication primitive algorithms, which embodies the concept of “Intra-Inter” co-design.
@@
“Host-Net” co-design primarily refers to the collaboration of computing resources between the end-host and programmable switches within a network. In recent years, with the development of programmable switch technology, in-network aggregation has gradually become a new computing paradigm. In-network aggregation uses the computing capabilities of programmable switches within the network to reduce the traffic running in the network. As shown by the blue line in Fig. 5(b), the two flows of Job2 are aggregated into one flow at the ToR switch, thus reducing further communication overhead. However, since existing protocols are designed for end-to-end communication, implementing in-network aggregation usually requires the support of new protocols. For instance, ATP [15] is an in-network aggregation transmission protocol designed for multi-tenant scenarios, which can make full use of the computing capabilities of multi-level programmable switches and degrade to host aggregation when the in-network computing resources are exhausted. This approach also reduces the network traffic for a collective communication task. ATP embodies the concept of “Host-Net” co-design. Nevertheless, in-network aggregation has not been widely applied in the production environment due to its complexity. “Host-Net” co-design remains a direction worthy of long-term research.
IV-C Open Issues
Cross-layer information exchange mechanism: The lack of effective information exchange mechanism across different layers makes collaborative optimization difficult. How to design good protocols to achieve efficient information exchange between different layers is a problem worth studying.
Utilization of heterogeneous network resources: The full utilization of heterogeneous network resources is a long-term complex problem, involving the awareness of the real-time status of different network resources and timely scheduling. This poses high requirements on the complexity and accuracy of the scheduler.
Dynamic optimization of communication libraries: Current communication optimization strategies are usually offline, and existing generative CCLs are difficult to meet the real-time requirements of online operation. However, network status is dynamic, and communication patterns also change in emerging model structures such as MoE. Therefore, the dynamic optimization of communication libraries is an important research direction.
Network topology specific for distributed training: Current network topologies are usually optimized for general tasks (e.g., fat-tree) or common communication primitive implementations (e.g., Ring-based All-Reduce). How to design or generate an efficient network topology for distributed training tasks and introduce dynamic adjustments to the topology according to the change in traffic load is a research direction worth exploring.
Communication Optimization with emerging techniques: With the emergence of new hardware (such as programmable switches, optical switches, etc.), it is important to study the impact of these new hardware on communication optimization and explore how to make full use of these hardware to improve communication efficiency.